Exercise 11-9 Mixed cost at different levels of activity LO 11-1

Zachary Corporation paid one of its sales representatives $8,000 during the month of March. The rep is paid a base salary plus $13 per unit of product sold. During March, the rep sold 100 units.


Required

Calculate the total monthly cost of the sales representative’s salary for each of the following months:

Month April May June July
Number of units sold 140 50 150 60
Total variable cost
Total fixed cost
Total salary cost
Answer #1
Month April May June July
Number of units sold 140 50 150 60
Total variable cost 1820 650 1950 780
Total fixed cost 6700 6700 6700 6700
Total salary cost 8520 7350 8650 7480
Workings:
Base salary 6700 =8000-(100*13)
April May June July
Total variable cost =140*13 =50*13 =150*13 =60*13